There are 83 students in school. The number of boys is 31 less than 2 times the number of girls. How many boys and girls are there

Answer by amarjeeth123(569)

Let the number of girls be x.
Then the number of boys is 2x-31.
Total number of students=83
x+2x-31=83
3x=83+31=114
x=38
Number of girls is 38.
Number of boys is 2(38)-31=76-31=45.
Number of girls is 38.
Number of boys is 45.